OpenStudy (anonymous):

solve using integrating factor y'=x-2y tnx:)

OpenStudy (anonymous):

y'+2y=x y'+P(x)y=Q(x) I=\(e^{\int P(x)dx}\) =\(e^{\int 2dx}\) = \(e^{2x}\) \[e^{2x}y'+2e^{2x}y=e^{2x}x\]\[\frac{d e^{2x} y}{dx} =e^{2x}x\]\[\int\frac{d e^{2x} y}{dx} =\int e^{2x}x dx\]\[e^{2x} y=\frac{1}{4}e^{2x}(2x-1)+C\]\[y=\frac{1}{4}(2x-1)+C\]